• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Olivier

    Log entry time 16:04:30 on April23,2001

    Entry number 61134

    keyword=Day shift summary

    8:00 Controlled access for set up of neutron detector test
    11:30 Close the hall, start OTR test (J.-C. Denard)
    12:10 JC Denard asks for 3 harp scan measurements
    12:20 JC is done with OTR test. Ask for access to check trigger cable on neutron detector.
    12:50 Start 3831 , LH2.
    Change prescale factors
    12:57 Start 3832
    13:21 Start 3833
    13:28 Start 3834, ped run
    13:36 Start 3835
    14:10 Controlled access for cabling work on neutron detector
    14:50 Start 3836
    15:42 Attempt to move to empty target. Motion fails. Reboot target IOC, decide to go to carbon target instead.

    Target motion problem not solved at 16:00. J.P called in.